OpenStudy (anonymous):

write the equation y= -6/5x-4 in standard form

OpenStudy (anonymous):

please explain

Nnesha (nnesha):

the standard form is \[\large\rm Ax+By=C\] where ` a` should be the positive integer

OpenStudy (anonymous):

but none are positive

Nnesha (nnesha):

alright then keep it negative

OpenStudy (anonymous):

how do i solve this?

Nnesha (nnesha):

well wait ..what do you mean ? `but none are positive`

OpenStudy (anonymous):

none of the numbers are positive

Nnesha (nnesha):

the standard form is Ax+By=C so we need to move the x term to the left side

Nnesha (nnesha):

how would you cancel out -6/5x from right side ?

OpenStudy (anonymous):

add -6/5x to both sides

Nnesha (nnesha):

hmmm it's a fraction for we need to get the common denominator are right side

OpenStudy (anonymous):

im sry can i talk later got busy

Nnesha (nnesha):

\[\huge\rm y=\frac{-6}{5}x-4\] multiply by the least common denominator of the fraction

Nnesha (nnesha):

\[\huge\rm y=\frac{-6}{\color{Red}{5}}x-4\] multiply both sides by 5 \[\huge\rm 5*y=5*(\frac{-6}{\color{Red}{5}}x-4)\] and then simiplify

Nnesha (nnesha):

don't forget to distribute parentheses by 5

OpenStudy (anonymous):

5*y=-1.2x-20

OpenStudy (anonymous):

5*y=-6x-20

OpenStudy (danjs):

yeah, just move the -6x over to get the standard form